3x+5=2x+2
4x+3=5x-4
2x-5=4x-1
5x+24=2x+15
5y-10=14-3y

Answer by CubeyThePenguin(3113)

3x + 5 = 2x + 2
x = -3

4x + 3 = 5x - 4
-x = -7
x = 7

2x - 5 = 4x - 1
-2x = 4
x = -2

5x + 24 = 2x + 15
3x = -9
x = -3

5y - 10 = 14 - 3y
8y = 24
y = 3